OpenStudy (anonymous):

Write an appropriate direct variation equation if y = 14 when x = -14. Answer y = 14x y = -14x y = x y = -x

jimthompson5910 (jim_thompson5910):

y = kx y = k*(-14) 14 = k*(-14) solve for k

OpenStudy (anonymous):

y=-x @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good, k = -1, so y = -1*x or y = -x